# Kinetic study of Kelvin-Helmholtz instability: Multiscale thermodynamic nonequilibrium effects and their relative importance

### v1

@@ Abstract
This study investigates the complex dynamics of thermodynamic nonequilibrium effects (TNEs) and their relative importance during the development of the Kelvin-Helmholtz instability (KHI) using high-order discrete Boltzmann models (DBMs). First, the capabilities and differences among various discrete velocity sets in capturing TNEs and distribution functions are assessed. Based on this analysis, practical guidelines for constructing discrete velocity stencils are proposed to enhance phase-space discretization and improve the robustness of high-order DBMs. At different stages of KHI and under varying initial conditions, multiscale TNEs, such as viscous stresses of different orders, emerge with distinct dominant roles. Specifically, three scenarios are identified: (i) regimes dominated by first-order TNEs, (ii) alternation between first- and second-order TNEs, and (iii) states where second-order TNEs govern the system’s behavior. To quantitatively capture these transitions, criteria for TNE dominance at different orders in KHI evolution are established based on the relative thermodynamic nonequilibrium intensity ($R_{\text{TNE}}$). In scenarios dominated by second-order TNEs, differences between first-order and second-order models are compared in terms of macroscopic quantities, nonequilibrium effects, and moment relations, revealing the physical limitations of low-order models in capturing TNEs. Furthermore, the effectiveness, extensibility, and limitations of a representative high-order model are examined under second-order TNE-dominated conditions. To encapsulate these findings, a nonequilibrium phase diagram that visually maps the multiscale characteristics of KHI is constructed. This diagram not only provides intuitive insights into the dynamic interplay of different nonequilibrium effects but also serves as a kinetic roadmap for selecting suitable models under diverse nonequilibrium conditions.
@@ I Introduction
In 2012, Xu et al. introduced the concept of utilizing non-conserved kinetic moments of the distribution function to characterize a system’s nonequilibrium state, effects, and behaviors, marking the inception of DBM researchXu et al. (2012). In 2024, Xu et al. further expanded the DBM framework, refining its coarse-grained modeling strategies and complex physical field analysis. They elucidated its connections to traditional fluid modeling and other kinetic methods, providing insights for the continuous progress of DBM Xu, Zhang, and Gan (2024).
@@
The current DBM adopts a traditional modeling and simulation approach, where physical principles and numerical computation techniques are considered as distinct processes. As both a kinetic model construction method and a tool for analyzing complex physical fields, DBM focuses on the following two fundamental aspects: (i) Model construction, which defines the governing equations of the model. The governing equations consist of the evolution equation for system behavior and the discrete velocity constraint equations. (ii) Analysis level: By leveraging higher-order non-conserved moments of $(f-f^{eq})$, DBM efficiently extracts discrete and nonequilibrium states, effects, and behaviors. Additionally, constructing the phase space via the independent components of $\mathbf{M}_{m,n}(f-f^{eq})$ provides an intuitive representation of these nonequilibrium characteristics. In particular, DBM resolves two critical questions: “how to describe discrete states and effects”, and “how to extract and analyze nonequilibrium states and effects”. The refinement of DBM modeling strategies has further clarified its physical applicability. In recent years, DBM has been extensively applied to a wide range of complex fluid systems, including: liquid-vapor phase-separating systems Gan et al. (2015, 2022), hydrodynamic instabilitiesLai et al. (2016); Chen et al. (2025, 2024); Xu, Lin, and Lai (2025), reactive flows (e.g., detonation and combustion) Zhang et al. (2016); Lin et al. (2016a); Ji, Lin, and Luo (2022), Shock reflection and shock wave/boundary layer interaction Bao et al. (2022); Qiu et al. (2020, 2021), micro-nanoscale flowsZhang et al. (2019, 2023), plasma kineticsLiu et al. (2023); Song et al. (2024), etc.
@@
In recent years, significant progress has been made in studying KHI and its nonequilibrium effects using DBM Gan et al. (2011); Lin et al. (2016b); Gan et al. (2019); Lin et al. (2019, 2021); Li and Lin (2024).
@@
Recent insights: Most recently, in 2024, Li et al. used a two-component DBM to study the effects of initial amplitude and Reynolds number ($Re$) on KHI in the presence of a gravitational field. Their results provided a comprehensive understanding of how density gradients, mixing degree, mixing width, viscous stress, and heat flux are influenced by these factors, offering new perspectives on KHI under complex initial and boundary conditions Li and Lin (2024).
@@ II Discrete Boltzmann modeling and complex physical field analysis method
As nonequilibrium intensity increases, the derivation and solution of the EHEs become significantly more challenging due to their pronounced nonlinearity and the presence of numerous nonlinear terms. In multiphase systems, where different media possess distinct velocities and mean flow velocities, the EHEs’ forms are not unique. In contrast, the complexity of DBM modeling increases only marginally under such conditions. This is achieved by incorporating the necessary elements into $\bm{\Psi}^{\prime}\left(\mathbf{v}_{i},\eta_{i}\right)$ Gan et al. (2022). Specifically, when the Knudsen number increases by an order of magnitude, DBM requires the introduction of only two additional conserved moments. This advantage renders DBM particularly well-suited for modeling and simulating nonequilibrium complex fluids Xu, Zhang, and Gan (2024).
@@ III Key considerations for constructing discrete velocity sets
As a theoretical framework for model construction and complex physical field analysis, the DBM imposes only the essential physical constraints required to describe the considered TNEs. It does not specify temporal, spatial, or phase-space discretization schemes. However, numerically, the choice of DVS plays a crucial role in determining the DBM’s ability to accurately capture physical effects. In 2018, Gan et al. proposed a two-dimensional discrete velocity set with 26 discrete velocities (D2V26) for phase-space discretization Gan et al. (2018). Building on this foundation, four new velocity sets (DVS1-DVS4) are developed to incorporate second-order nonequilibrium effects while optimizing computational efficiency and stability (Figure 1). The key characteristics of these sets are as follows: (i) DVS1: Retains the first 25 discrete velocities from Ref. Gan et al., 2018. (ii) DVS2: A more symmetric 25-velocity set designed for improved performance. (iii) DVS3: Extends DVS2 by introducing additional velocity magnitudes, resulting in a more diverse and evenly distributed velocity set. (iv) DVS4: Further optimizes velocity distribution by increasing the diversity of magnitude combinations. The detailed configurations of the four DVSs are given as follows:
@@
In contrast, the simulation results obtained using DVS4 closely match the second-order analytical solution for viscous stress while demonstrating excellent numerical stability. These findings underscore that a well-designed DVS is fundamental to the physical fidelity of DBM, ensuring both stability and accuracy in modeling nonequilibrium effects.
@@
Numerical stability is a critical concern in computational fluid dynamics, particularly in nonequilibrium simulations. The primary factors influencing numerical stability can be broadly classified into (i) the numerical schemes for spatiotemporal derivative computations and (ii) the discretization of phase space. Regarding numerical schemes, all four DVSs employ a fifth-order weighted essentially non-oscillatory finite difference scheme for spatial derivatives and a second-order implicit Runge-Kutta finite difference scheme for temporal derivatives. These high-order schemes are sufficiently robust to suppress numerical oscillationsJiang and Shu (1996); Ascher, Ruuth, and Spiteri (1997). Consequently, the observed oscillations primarily stem from the discretization of phase space, which plays a crucial role in preserving moment relations post-discretization. This, in turn, influences the effectiveness of the various DVSs in implementing the functionality of the original physical model.
In fact, the preservation of moment relations can be further categorized into theoretical preservation and numerical preservation. Theoretical preservation ensures the existence of the inverse matrix for the DVS when solving the equilibrium distribution function. This condition is satisfied by all the aforementioned DVSs. Numerical preservation, however, is assessed by examining the smoothness of the distribution function $f$ during the simulation process. Next, we analyze the ability of the four DVSs to describe the distribution function $f$ and assess their impact on numerical stability and accuracy.
The preservation of moment relations can be further categorized into two aspects: theoretical preservation and numerical preservation. Theoretical preservation guarantees the existence of an invertible matrix for the DVS, ensuring a well-defined equilibrium distribution function. This condition is met by all the aforementioned DVSs. Numerical preservation, however, is evaluated based on the smoothness of the distribution function $f$ throughout the simulation. Figure 3 presents distributions of $f$ and $\left(f-f^{eq}\right)$ along the horizontal line $y=0.3L_{y}$ at $t=0.017$, as computed using the four DVSs. Key observations are as follows:
@@
Based on this analysis, DVS4 is selected for subsequent numerical simulations. This study offers practical insights into optimizing phase space discretization and developing robust, high-order DBM models.
@@ IV Multiscale TNEs and their relative importance in KHI
(ii) Comparison of high-order and low-order models: When higher-order nonequilibrium effects become significant, the differences between high-order and low-order models are evaluated. The analysis focuses on macroscopic quantities, nonequilibrium effects, and moment relationships, highlighting the limitations of low-order models in capturing complex nonequilibrium dynamics.
@@
The perturbation amplitude is set to $u_{0}=50\Delta y$, and the initial perturbation wavenumber is given by $k={2\pi}/{L_{y}}$. Boundary conditions for the KHI simulations are implemented as follows: periodic boundary conditions are applied in the $y$-direction, while outflow boundary conditions, commonly used in KHI studies, are enforced in the $x$-direction Keppens and Tóth (1999); Perucho et al. (2007); Lifeng et al. (2009).
@@ IV.2.1 Comparison of viscous stress and its components
Next, we analyze the physical mechanisms governing the onset of second-order nonequilibrium effects by decomposing the analytical expression of $\Delta_{2xx}^{*(2)}$ into its fundamental components. Based on the nature of the driving forces, second-order nonequilibrium effects can be decomposed into four terms. Velocity gradient term:
@@
Temperature gradient term:
@@
Second-order density gradient term:
@@
First-order density gradient term:
@@ IV.2.2 Comparison of KHI growth rate
(ii) In the D2V16 DVS, $\ln\left(\left.E_{x}\right|_{\max}\right)$ diverges for $t>0.8$, demonstrating the model’s failure to capture viscous stress accurately. Throughout KHI evolution, the D2V16 model substantially overestimates TNEs, leading to simulation divergence.
@@
Kinetic moment relations are fundamental physical quantities that characterize the nonequilibrium state and dynamics of fluid systems. Increasing the number and order of preserved moment relations in physical modeling enhances the model’s ability to capture the fluid’s nonlinear dynamics. To directly assess the limitations of the low-order model and the advantages of the high-order model, we compare the moment relations computed by both approaches.
@@
The D2V25 DVS preserves 25 independent moment relations, whereas the D2V16 DVS retains only the first 16. Figure 12 shows the kinetic moment relations $\mathbf{M}_{0},\mathbf{M}_{1},\mathbf{M}_{2,0},\mathbf{M}_{2},\mathbf{M}_{3,1},\mathbf{M}_{3}$, and $\mathbf{M}_{4,2}$ along $y=0.7L_{y}$ at $t=1.68$ for both D2V25 and D2V16 DVSs, along with their differences. The relative error $\varepsilon={|M_{\alpha\beta 25}-M_{\alpha\beta 16}|}/{|M_{\alpha\beta 25}|}$ in the D2V16 model accumulates as KHI evolves. As shown in Fig. 7, at $t=1.68$, large-scale vortices form along the fluid interface. At this stage, the accumulated relative errors in the first 16 moment relations of D2V16 become significant. Specifically, in the distribution curves of $M_{1y}$, $M_{2xy}$, $M_{3,1y}$, $M_{3xxy}$, and $M_{4,2xy}$, the maximum relative errors of D2V16 compared to D2V25 are as follows: 200% at $x=130$ for $M_{1y}$, 110% at $x=85$ for $M_{2xy}$, 160% at $x=110$ for $M_{3,1y}$, 100% at $x=105$ for $M_{3xxy}$, 240% at $x=105$ for $M_{4,2xy}$. Additionally, several other moment relations exhibit errors exceeding $50\%$, further highlighting the limitations of the low-order model.
@@
The relative errors in the first 16 moment relations of the D2V16 model primarily arise from its failure to satisfy higher-order moment conditions, particularly $\mathbf{M}_{4}$ and $\mathbf{M}_{5,3}$. Satisfying the first 16 moment relations ensures the recovery of linear constitutive laws, while higher-order moment conditions govern nonlinear constitutive behavior. As the relaxation time increases and shear velocity intensifies, TNEs exhibit strong nonlinearity. Consequently, the linear constitutive laws in the D2V16 model lead to significant macroscopic errors. In contrast, the D2V25 DVS satisfies all the moment conditions necessary to accurately capture higher-order constitutive effects.
@@
Figure 13 presents the calculations of the moment relations $\mathbf{M}_{4}$ and $\mathbf{M}_{5,3}$ along the line $y=0.7L_{y}$ at $t=1.68$, obtained using the D2V25 and D2V16 models, along with their differences. The description of higher-order moment relations by the low-order model is essentially a “guess”, which proves highly inaccurate. As shown in Fig. 13, the “guess” made by the D2V16 model is highly inaccurate. For $\mathbf{M}_{4}$, the relative error in the D2V16 model reaches: 300% at $x=110$ for $M_{4xxxx}$ and $M_{4yyyy}$, 340% at $x=110$ for $M_{4xxyy}$. Additionally, several other moment relations exhibit relative errors of approximately 100%. This further highlights the limitations of the low-order model in capturing higher-order nonequilibrium effects.
IV.2.3 Comparison of Macroscopic Quantities

### v2

# Multiscale thermodynamic nonequilibrium effects in Kelvin-Helmholtz instability and their relative importance

@@ Abstract
This study investigates the complex kinetics of thermodynamic nonequilibrium effects (TNEs) and their relative importance during the development of the Kelvin-Helmholtz instability (KHI) using high-order discrete Boltzmann models (DBMs). First, the capabilities and differences among various discrete velocity sets in capturing TNEs and distribution functions are assessed. Based on this analysis, practical guidelines for constructing discrete velocity stencils are proposed to enhance phase-space discretization and improve the robustness of high-order DBM simulation. At different stages of KHI and under varying initial conditions, multiscale TNEs, such as viscous stresses of different orders, emerge with distinct dominant roles. Specifically, three scenarios are identified: (i) regimes dominated by first-order TNEs, (ii) alternation between first- and second-order TNEs, and (iii) states where second-order TNEs govern the system’s behavior. To quantitatively capture these transitions, criteria for TNE dominance at different orders in KHI evolution are established based on the relative thermodynamic nonequilibrium intensity ($R_{\text{TNE}}$). In scenarios dominated by second-order TNEs, differences between first-order and second-order models are compared in terms of macroscopic quantities, nonequilibrium effects, and kinetic moments, revealing the physical limitations of low-order models in capturing TNEs. Furthermore, the effectiveness, extensibility, and limitations of a representative high-order model are examined under second-order TNE-dominated conditions. To encapsulate these findings, a nonequilibrium phase diagram that visually maps the multiscale characteristics of KHI is constructed. This diagram not only provides intuitive insights into the dynamic interplay of different nonequilibrium effects but also serves as a kinetic roadmap for selecting suitable models under diverse nonequilibrium conditions.
@@ I Introduction
DBM evolved from the physical modeling branch of LBM. Before 2012, the two methods had no differences in physical functionality. In 2012, Xu et al. introduced the use of non-conserved kinetic moments of the distribution function to characterize a system’s nonequilibrium state, effects, and behaviors. This marked the inception of DBM research and its distinction from standard LBM in terms of physical functionalityXu et al. (2012). In 2024, Xu et al. further expanded the DBM framework, refining its coarse-grained modeling strategies and complex physical field analysis. They elucidated its connections to traditional fluid modeling and other kinetic methods, providing insights for the continuous progress of DBM Xu, Zhang, and Gan (2024).
@@
Currently, we are following the traditional research route where the physics model construction and the numerical scheme for solving model equations are investigated separately. As both a kinetic model construction method and a tool for analyzing complex physical fields, DBM focuses on two fundamental aspects: (i) Model construction: This defines the governing equations, which include the evolution equation governing system behavior and the discrete velocity constraint equations. (ii) Analysis framework: By leveraging higher-order non-conserved moments of $(f-f^{eq})$, DBM efficiently extracts discrete and nonequilibrium states, effects, and behaviors. Additionally, constructing the phase space via the independent components of $\mathbf{M}_{m,n}(f-f^{eq})$ provides an intuitive representation of these nonequilibrium characteristics.
In particular, DBM addresses two critical questions: “how to include sufficient discrete/TNE states and effects in the modeling equations”, and “how to extract and analyze the underlying discrete/TNE states and effects from the massive simulation data”. The former is point beyond the NS modeling. The latter is the point beyond current other kinetic simulation methods. The refinement of DBM strategies has further clarified its physical applicability.
In recent years, DBM has been extensively applied to a wide range of complex fluid systems, including: liquid-vapor phase-separating systems Gan et al. (2015, 2022), hydrodynamic instabilitiesLai et al. (2016); Chen et al. (2025, 2024); Xu, Lin, and Lai (2025), reactive flows (e.g., detonation and combustion) Zhang et al. (2016); Lin et al. (2016a); Ji, Lin, and Luo (2022), Shock reflection and shock wave/boundary layer interaction Bao et al. (2022); Qiu et al. (2020, 2021), micro-nanoscale flowsZhang et al. (2019, 2023), plasma kineticsLiu et al. (2023); Song et al. (2024), etc. As an analytical tool for complex physical fields, DBM can be used to characterize the evolution stages of complex flows and analyze the physical mechanisms of flow systems. For instance, the temporal evolution of non-equilibrium intensity $D$ has been employed to distinguish different stages of phase separation, including the spinodal decomposition stage, the transition to domain growth, and the domain growth stage Gan et al. (2015). Similarly, the evolution of the non-equilibrium component $\Delta_{3,1y}^{*}$ has been used to characterize various stages of RTI development and serve as a physical marker for interface tracking in complex flow simulations Lai et al. (2016). The evolution of viscous stress has been utilized to classify different stages of bubble coalescence Sun et al. (2024), etc.
@@
In recent years, significant progress has also been made in studying KHI and its nonequilibrium effects using DBM Gan et al. (2011); Lin et al. (2016b); Gan et al. (2019); Lin et al. (2019, 2021); Li and Lin (2024).
@@
Recent insights: Most recently, in 2024, Li et al. used a two-component DBM to study the effects of initial amplitude and $Re$ on KHI in the presence of a gravitational field. Their results provided a comprehensive understanding of how density gradients, mixing degree, mixing width, viscous stress, and heat flux are influenced by these factors, offering new perspectives on KHI under complex initial and boundary conditions Li and Lin (2024).
@@ II Discrete Boltzmann modeling and complex physical field analysis method
As nonequilibrium intensity increases, the derivation and solution of the EHEs become significantly more challenging due to their pronounced nonlinearity and the presence of huge amount of nonlinear terms. In multiphase systems, where different media possess distinct velocities and mean flow velocities, the EHEs’ forms are not unique. In contrast, the complexity of DBM modeling increases only marginally under such conditions. This is achieved by incorporating the necessary elements into $\bm{\Psi}^{\prime}\left(\mathbf{v}_{i},\eta_{i}\right)$ Gan et al. (2022). Specifically, when the Knudsen number ($Kn$) increases by an order of magnitude, DBM requires the introduction of only two additional non-conserved moments. This advantage renders DBM particularly well-suited for modeling and simulating nonequilibrium complex fluids Xu, Zhang, and Gan (2024).
@@ III Key considerations for constructing discrete velocity sets
As a theoretical framework for model construction and complex physical field analysis, the DBM imposes only the essential physical constraints required to describe the considered TNEs. It does not specify temporal, spatial, or phase-space discretization schemes. However, numerically, the choice of DVS plays a crucial role in determining the ability to accurately capture physical effects in DBM simulation. In 2018, Gan et al. proposed a two-dimensional discrete velocity set with 26 discrete velocities (D2V26) for phase-space discretization Gan et al. (2018). Building on this foundation, four new velocity sets (DVS1-DVS4) are developed to incorporate second-order nonequilibrium effects while optimizing computational efficiency and stability (Figure 1). The key characteristics of these sets are as follows: (i) DVS1: Retains the first 25 discrete velocities from Ref. Gan et al., 2018. (ii) DVS2: A more symmetric 25-velocity set designed for improved performance. (iii) DVS3: Extends DVS2 by introducing additional velocity magnitudes, resulting in a more diverse and evenly distributed velocity set. (iv) DVS4: Further optimizes velocity distribution by increasing the diversity of magnitude combinations. The detailed configurations of the four DVSs are given as follows:
@@
In contrast, the simulation results obtained using DVS4 closely match the second-order analytical solution for viscous stress while demonstrating excellent numerical stability. These findings underscore that a well-designed DVS is fundamental to the physical fidelity of DBM simulation, ensuring both stability and accuracy in modeling nonequilibrium effects.
@@
Numerical stability is a critical concern in computational fluid dynamics, particularly in nonequilibrium simulations. The primary factors influencing numerical stability can be broadly classified into (i) the accuracy of physical model, (ii) the numerical schemes for spatiotemporal derivative computations and (iii) the discretization of phase space. Regarding the accuracy of physical models, all four DVSs belong to high-order models that surpass the NS level. High-order DBM integrates high-order constitutive relations to describe high-order nonequilibrium effects, enabling a more precise capture of small-scale structures and dynamical patterns, as well as their interactions. These models not only enhance physical functionality but also improve numerical stability by accurately representing the intrinsic driving mechanisms of system evolution. Regarding numerical schemes, all four DVSs employ a fifth-order weighted essentially non-oscillatory finite difference scheme for spatial derivatives and a second-order implicit Runge-Kutta finite difference scheme for temporal derivatives. These high-order schemes are sufficiently robust to suppress numerical oscillationsJiang and Shu (1996); Ascher, Ruuth, and Spiteri (1997). Consequently, the observed oscillations primarily stem from the discretization of phase space, which plays a crucial role in preserving moment relations post-discretization. This, in turn, influences the effectiveness of the various DVSs in implementing the functionality of the original physical model.
The preservation of kinetic moments can be further categorized into two aspects: theoretical preservation and numerical preservation. Theoretical preservation guarantees the existence of an invertible matrix for the DVS, ensuring a well-defined equilibrium distribution function. This condition is met by all the aforementioned DVSs. Numerical preservation, however, is evaluated based on the smoothness of the distribution function $f$ throughout the simulation. Figure 3 presents distributions of $f$ and $\left(f-f^{eq}\right)$ along the horizontal line $y=0.3L_{y}$ at $t=0.017$, as computed using the four DVSs. Key observations are as follows:
@@
Based on this analysis, DVS4 is selected for subsequent numerical simulations. This study provides practical insights into optimizing phase space discretization and developing robust, high-order DBM simulation. In future work, we will apply von Neumann stability analysis to theoretically investigate its impact on numerical stability Gan et al. (2008).
@@ IV Multiscale TNEs and their relative importance in KHI
(ii) Comparison of high-order and low-order models: When higher-order nonequilibrium effects become significant, the differences between high-order and low-order models are evaluated. The analysis focuses on macroscopic quantities, nonequilibrium effects, and kinetic moments, highlighting the limitations of low-order models in capturing complex nonequilibrium dynamics.
@@
The perturbation amplitude is set to $u_{0}=50\Delta y$, and the initial perturbation wavenumber is given by $k={2\pi}/{L_{y}}$.
Boundary conditions for the KHI simulations are implemented as follows: periodic boundary conditions are applied in the $y$-direction, while outflow boundary conditions, commonly used in KHI studies, are enforced in the $x$-direction Keppens and Tóth (1999); Perucho et al. (2007); Lifeng et al. (2009). Boundary conditions play a critical role in fluid models, directly influencing nonequilibrium behaviors. Properly designed kinetic boundary conditions are essential for ensuring physical accuracy, numerical stability, and computational precision.
@@ IV.2.1 Comparison of viscous stress and its components
Next, we analyze the physical mechanisms governing the onset of second-order nonequilibrium effects by decomposing the analytical expression of $\Delta_{2xx}^{*(2)}$ into its fundamental components. Based on the nature of the driving forces, second-order nonequilibrium effects can be decomposed into four termsGan et al. (2025).
(i) The velocity gradient term $T_{u}$ is composed of the squares or products of velocity gradients, capturing the nonlinear effects of velocity variations:
@@
(ii) The temperature gradient term $T_{t}$ is related to the square of the temperature gradient, reflecting the intensity of temperature gradients within the system:
@@
(iii) The second-order density gradient term $T_{\rho 1}$ involves the second derivatives of density, reflecting the density curvature in the $x$- and $y$-directions:
@@
(iv) The first-order density gradient term $T_{\rho 2}$ is related to the square of the first-order density gradient, reflecting the role of density inhomogeneity in viscous stress:
@@ IV.2.2 Comparison of KHI growth rate
(ii) In the D2V16 DVS, $\ln\left(\left.E_{x}\right|_{\max}\right)$ diverges for $t>0.8$, demonstrating the model’s failure to capture viscous stress accurately. Throughout KHI evolution, the D2V16 DVS substantially overestimates TNEs, leading to simulation divergence.
@@ IV.2.3 Comparison of kinetic moments
Kinetic moments are fundamental physical quantities that characterize the nonequilibrium state and dynamics of fluid systems. Increasing the number and order of preserved kinetic moments in physical modeling enhances the model’s ability to capture the fluid’s nonlinear dynamics. To directly assess the limitations of the low-order model and the advantages of the high-order model, we compare the kinetic moments computed by both approaches.
@@
The D2V25 DVS preserves 25 independent kinetic moments, whereas the D2V16 DVS retains only the first 16. Figure 12 shows the kinetic moments $\mathbf{M}_{0},\mathbf{M}_{1},\mathbf{M}_{2,0},\mathbf{M}_{2},\mathbf{M}_{3,1},\mathbf{M}_{3}$, and $\mathbf{M}_{4,2}$ along $y=0.7L_{y}$ at $t=1.68$ for both D2V25 and D2V16 DVSs, along with their differences. The relative error $\varepsilon={|M_{\alpha\beta 25}-M_{\alpha\beta 16}|}/{|M_{\alpha\beta 25}|}$ in the D2V16 DVS accumulates as KHI evolves. As shown in Fig. 7, at $t=1.68$, large-scale vortices form along the fluid interface. At this stage, the accumulated relative errors in the first 16 kinetic moments of D2V16 become significant. Specifically, in the distribution curves of $M_{1y}$, $M_{2xy}$, $M_{3,1y}$, $M_{3xxy}$, and $M_{4,2xy}$, the maximum relative errors of D2V16 compared to D2V25 are as follows: 200% at $x=130$ for $M_{1y}$, 110% at $x=85$ for $M_{2xy}$, 160% at $x=110$ for $M_{3,1y}$, 100% at $x=105$ for $M_{3xxy}$, 240% at $x=105$ for $M_{4,2xy}$. Additionally, several other kinetic moments exhibit errors exceeding $50\%$, further highlighting the limitations of the low-order model.
@@
The relative errors in the first 16 kinetic moments of the D2V16 DVS primarily arise from its failure to satisfy higher-order moment conditions, particularly $\mathbf{M}_{4}$ and $\mathbf{M}_{5,3}$. Satisfying the first 16 kinetic moments ensures the recovery of linear constitutive laws, while higher-order moment conditions govern nonlinear constitutive behavior. As the relaxation time increases and shear velocity intensifies, TNEs exhibit strong nonlinearity. Consequently, the linear constitutive laws in the D2V16 DVS lead to significant macroscopic errors. In contrast, the D2V25 DVS satisfies all the moment conditions necessary to accurately capture higher-order constitutive effects.
@@
Figure 13 presents the calculations of the kinetic moments $\mathbf{M}_{4}$ and $\mathbf{M}_{5,3}$ along the line $y=0.7L_{y}$ at $t=1.68$, obtained using the D2V25 and D2V16 DVSs, along with their differences. The description of higher-order kinetic moments by the low-order model is essentially a “guess”, which proves highly inaccurate. As shown in Fig. 13, the “guess” made by the D2V16 DVS is highly inaccurate. For $\mathbf{M}_{4}$, the relative error in the D2V16 DVS reaches: 300% at $x=110$ for $M_{4xxxx}$ and $M_{4yyyy}$, 340% at $x=110$ for $M_{4xxyy}$. Additionally, several other kinetic moments exhibit relative errors of approximately 100%. This further highlights the limitations of the low-order model in capturing higher-order nonequilibrium effects.
IV.2.4 Comparison of macroscopic quantities